OpenStudy (anonymous):

How would I answer this question? 16^(x-1)=8

OpenStudy (anonymous):

\[16^{x-1}=8\]

OpenStudy (anonymous):

take logs. (x - 1) ln 16 = ln 8 x - 1 = (ln 8) / (ln 16) x = (ln 8) / (ln 16) + 1 = 1.75

OpenStudy (anonymous):

All good now, @lunatic09 ?

OpenStudy (anonymous):

Good luck to you in all of your studies! @lunatic09

OpenStudy (anonymous):

Thank you so much, I get it!

OpenStudy (anonymous):

Im gona have to look up log rules, aghhh haha. Thanks again!

OpenStudy (anonymous):

uw! There is another "intuitive" way. You could look for a common root for a power and see "2". You would then take the 1/4 root of 16 to get "2" and then cube to get 8. That's a sort of short-cut.

OpenStudy (anonymous):

Oh I see 17^.75 is the same thing as 2^3. My brain is so rusty after a month of school :(, thanks for the shortcut.

OpenStudy (anonymous):

\[16^{x-1}=8\] can also be written as \[2^{4(x-1)}= 2^{3}\] since the bases are the same (2), equate the powers, that is 4(x-1)=3 4x-4=3 4x=3+4 4x=7 Hence, x= 7/4 or 1.75

OpenStudy (anonymous):

Thanks for the other method sisi_m

OpenStudy (anonymous):

So, [16^(1/4)]^3 -> 16^(3/4) = 8

OpenStudy (anonymous):

And since the exponent is "x - 1" we have: x - 1 = 3/4 -> x = 1.75 After a while, you'll do those in your head. Or go crazy first trying to!

OpenStudy (anonymous):

My brain is definitely going crazy right now, but it time Ill definitely get better. Thanks for the shortcut!

OpenStudy (anonymous):

So, that's actually 3 ways to do the problem. My 2 and also the method from sisi. Have a great day!

OpenStudy (anonymous):

You're a life saver, you too!